9. copts Find the taylor series of fl) se e-2x with Xo=0 use any method! show work
Answer #1

A taylor x=a ? by taylor series at given f(x) = 8 h (a) (x-a) kao k! 0 In our case, K +6w) = Ś f%) x ko KI *w = 10)+ flo) x
-22 16 e - 16 2개 F(x) = ( + C-2) + (4) 2. (-3)3 , 16x+ 32) C), tr(s) - Crear), the (0) (x) (Iker) -32e f() = -32e -32 F(x) =